Is the Function f(x)=|x²-x| differentiable at x=2. If yes find it`s derivative?

Grade:11

1 Answers

Sher Mohammad IIT Delhi
10 years ago
yes,

for x>1 f(x)=x^2-x
f'(2)=2*2-1=3

Sher Mohammad
B.Tech, IIT Delhi